An electron is to be accelerated from a velocity of 2.5x10^6 m/s to a velocity of 8x10^6 m/s.
A. Through what potential difference must the electron pass to accomplish this ?
B. Through what potential difference must the electron pass if it is to be slowed from 8x10^6 m/s to a halt?
Answer #1

Solution:

A) Work done = change in K.E

=> W = 0.5 * m ( (8*10^6)^2 - (2.5*10^6)^2 )

=> W = 0.5 * 9.1*10^-31 * 57.75 * 10^12

=> W = 262.7625 * 10^-19 V

=> V * q = 262.7625 * 10^-19 V

=> V * 1.6 * 10^-19 = 262.7625 * 10^-19

=> V = 164.2 Volt

b) Similarly V * q = 0.5 * 9.1 * 10^-31 * 64 * 10^12

=> V * 1.6 * 10^-19 = 291.2 * 10^-19

=> V = 182 V

please comment for Queries.

Please rate thanks.